# Size Generalization of Graph Neural Networks on Biological Data: Insights and Practices from the Spectral Perspective

### v4

@@ Abstract
noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt
1 Introduction
Graph neural networks (GNNs) (Kipf & Welling, 2017; Veličković et al., 2018; Hamilton et al., 2017; Zhang et al., 2018; Yan et al., 2019; Lee et al., 2018) have gained widespread popularity in graph classification tasks owing to their outstanding performance. Though most GNNs can process graphs of varying sizes, it remains under-explored whether they can generalize to graphs larger than those encountered during training (size generalizability). Size generalization in GNNs holds significant importance across multiple domains. For instance, in graph algorithmic reasoning (Yan et al., 2020; Veličković et al., 2020), GNNs are expected to learn complex algorithms from small examples and generalize that reasoning to larger graphs, as obtaining exact solutions for larger graphs is challenging. In the realm of biology, datasets exhibit a wide range of graph sizes, spanning from small molecules to large compounds. Evaluating whether learned knowledge is influenced by graph size is crucial, as size-dependent information may potentially have a detrimental impact on performance when employing pre-training strategies (Hu et al., 2020c).
Existing literature presents conflicting conclusions on GNNs’ size generalizability. On one hand, several studies (Levie et al., 2021; Xu et al., 2021; Sanchez-Gonzalez et al., 2020) have provided support for the ability of GNNs to effectively generalize across varying sizes. For instance, a theoretical study (Levie et al., 2021) established that spectral GNNs exhibit robust transferability between graphs with different sizes and topologies, provided that these graphs discretize the same underlying space in some generic sense. Other works further provided empirical evidence supporting the strong size generalizability of GNNs in the domains of algorithmic task learning (Xu et al., 2021) and physics simulations (Sanchez-Gonzalez et al., 2020). On the other hand, several studies  (Yehudai et al., 2021; Buffelli et al., 2022; Bevilacqua et al., 2021) have observed performance degradation when a size shift exists between the training and test data. For instance, a recent work (Yehudai et al., 2021) showed theoretically and empirically that the difference in d-patterns between small and large graphs contributes to this performance decline. However, we find that d-patterns cannot effectively capture variations in cycle statistics (Appendix K), a critical factor for the generalization of GNNs to larger biological graphs, as highlighted in this paper. There have also been proposals of novel models (Bevilacqua et al., 2021) and regularization techniques (Buffelli et al., 2022) to enhance the size generalizability of GNNs. These conflicts mainly arise from disparities in application domains and underlying assumptions concerning size-induced distribution shifts.
Motivated by these conflicts, we take a data-driven approach: we focus on real biological datasets and seek to characterize the types of size-induced distribution shifts. This characterization provides valuable insights into the size generalizability of GNNs. Specifically, we adopt a spectral perspective and identify the connections between the spectrum differences induced by varying graph sizes and the differences in subgraph patterns, particularly cycles. We find that breaking cycles in graphs amplifies the spectrum difference between smaller and larger graphs, whereas extending cycle lengths in smaller graphs to align with those in larger graphs reduces this difference. Furthermore, we observe that conventional GNNs struggle to generalize effectively without explicit cycle information, leading to performance degradation on larger graphs. While prior research has emphasized GNNs’ limitations in counting and recognizing cycles (Chen et al., 2020) within in-distribution generalization, this paper systematically explores their impact in out-of-distribution scenarios, specifically focusing on size generalizability. To address the above issues, we introduce a Size-Insensitive Attention (SIA) approach to enhance any GNNs. This is achieved by encoding cycle information into the feature representations and leveraging them within a size-insensitive attention mechanism to guide the learning process. To substantiate the utility of cycle information, we examine the performance of two additional simple baselines that also incorporate cycle information. Empirical evidence confirms that integrating cycle information indeed aids in GNN size generalization, and our proposed SIA strategy substantially boosts model performance on large test sets and surpasses the state-of-the-art models by a large margin.
In sum, our paper makes the following contributions:
  • New Observations. We characterize the types of distribution shifts caused by various graph sizes in biological networks, offering insights for designing a size-agnostic GNN.
  • Spectral Analysis. Unlike prior work, we leverage spectral analysis to deepen our understanding of the size generalizability of GNNs.
  • Effective Model-agnostic Strategy. To make GNNs aware of important size-related subgraph patterns (e.g., average cycle lengths), we propose a size-insensitive attention that can seamlessly integrate with various architectures. Experimental findings indicate that our method significantly enhances the size generalizability of GNNs and outperforms the baselines by a large margin.
2 Notations and Preliminaries
2.1 Notations & Definitions
Let ${\mathcal{G}}$(${\mathcal{V}}$, ${\mathcal{E}})$ be an undirected and unweighted graph with $N$ nodes, where ${\mathcal{V}}$ denotes the node set, and ${\mathcal{E}}$ denotes the edge set. The neighborhood of a node $v_{i}$ is defined as the set of all nodes that connect to $v_{i}$: ${\mathcal{N}}_{i}=\{v_{j}|(v_{j},v_{i})\in{\mathcal{E}}\}.$ The graph is represented by its adjacency matrix ${\mathbf{A}}\in\mathbb{R}^{N\times N}$, and it has a degree matrix ${\mathbf{D}}$, where the $i$th diagonal element $d_{i}$ corresponds to the degree of node $v_{i}$.
2.2 Graph Learning Task
In this paper, we focus on the graph classification task, where each node $v_{i}$ is associated with a feature vector ${\mathbf{x}}_{i}^{(0)}$, and the feature matrix ${\mathbf{X}}^{(0)}$ is constructed by arranging the node feature vectors as rows. When using a GNN for the graph classification task, we further denote the node representation matrix at the $l$-th layer as ${\mathbf{X}}^{(l)}$, and the representation of node $v_{i}$ as ${\mathbf{x}}_{i}^{(l)}$.
Supervised Graph Classification. Each graph ${\mathcal{G}}_{i}$ is associated with a label $y^{{\mathcal{G}}}_{i}$ sampled from a label set $\hat{{\mathcal{L}}}$. Given a subset of labeled graphs (from a label set $\hat{{\mathcal{L}}}$), the goal is to learn a mapping $f^{{\mathcal{G}}}:({\mathbf{A}},{\mathbf{X}}^{(0)})_{i}\mapsto y^{{\mathcal{G}}}_{i}$ between each graph ${\mathcal{G}}_{i}$ and its ground truth label $y^{{\mathcal{G}}}_{i}\in\hat{{\mathcal{L}}}$. The graph classification loss is given by $L$ = $\frac{1}{|{\mathcal{G}}_{\text{train}}|}\sum_{{\mathcal{G}}_{i}\in{\mathcal{G}}_{\text{train}}}$rossEntropy (${\mathbf{x}}^{{\mathcal{G}}_{i}}$, $y^{{\mathcal{G}}}_{i}$), where ${\mathcal{G}}_{\text{train}}$ is the training graph set and ${\mathbf{x}}^{{\mathcal{G}}_{i}}$ is the representation of graph ${\mathcal{G}}_{i}$.
Evaluation of Size Generalizability. Following prior work (Buffelli et al., 2022; Yehudai et al., 2021), we evaluate the size generalizability of GNNs by testing their classification performance on graphs whose sizes are larger than those in the train set. We obtain the small training graphs and large test graphs from the same dataset.
2.3 Graph Neural Networks
GNNs can be designed from either the spatial perspective or the spectral perspective. Despite the difference in the design perspectives, a recent work (Balcilar et al., 2021a) has shown that spectral GNNs and spatial GNNs are related and that spectral analysis of GNNs’ behavior can provide a complementary point of view to understand GNNs in general. Most spatial GNNs (Kipf & Welling, 2017; Xu et al., 2018; Veličković et al., 2018; Hamilton et al., 2017) use the message passing framework (Gilmer et al., 2017), which consists of three steps: neighborhood propagation, message combination and global pooling. Spectral GNNs (Bo et al., 2021; Defferrard et al., 2016; Levie et al., 2018) utilize the spectral properties of a propagation matrix ${\mathbf{T}}$ to perform the graph classification. The matrix ${\mathbf{T}}$ is usually a function of the adjacency matrix ${\mathbf{A}}$, such as the normalized adjacency matrix ${\mathbf{T}}=({\mathbf{D}}+{\mathbf{I}})^{-1/2}({\mathbf{A}}+{\mathbf{I}})({\mathbf{D}}+{\mathbf{I}})^{-1/2}$, or the normalized graph Laplacian matrix $\hat{{\mathbf{L}}}$. Since we consider an undirected graph with a real and symmetric adjacency matrix, the matrix ${\mathbf{T}}$ is also real and symmetric. Then, we can perform the eigendecomposition on the propagation matrix ${\mathbf{T}}$: ${\mathbf{T}}={\mathbf{U}}{\mathbf{\Lambda}}{\mathbf{U}}^{T}$, where ${\mathbf{U}}$ is an orthogonal matrix whose columns ${\mathbf{U}}_{i}$ are orthonormal and are the eigenvectors of ${\mathbf{T}}$, and ${\mathbf{\Lambda}}$ is a matrix whose diagonal elements are the eigenvalues of ${\mathbf{T}}$, sorted from large to small by their absolute values. The set of eigenvectors $\{{\mathbf{U}}_{i}\}$ form the orthonormal basis of ${\mathbb{R}}^{n}$. The goal of a spectral GNN is to learn a proper spectral filter: $f({\mathbf{\Lambda}})=c_{0}{\mathbf{I}}+c_{1}{\mathbf{\Lambda}}+c_{2}{\mathbf{\Lambda}}^{2}+\cdots+c_{i}{\mathbf{\Lambda}}^{i}+\cdots$, where $c_{i}$ are the learnable coefficients. The convolution at each layer can be viewed as or is equivalent to: ${\mathbf{X}}^{(l+1)}=\sigma({\mathbf{U}}f({\mathbf{\Lambda}}){\mathbf{U}}^{T}{\mathbf{X}}^{(l)}{\mathbf{W}}^{(l)})$, where ${\mathbf{W}}^{(l)}$ is a learnable weight matrix, and $\sigma{(\cdot)}$ is a nonlinear function (e.g., ReLU). The graph representation is obtained from the node representations at the last convolution layer: ${\mathbf{x}}^{{\mathcal{G}}}$ = Pooling ($\{{\mathbf{x}}_{i}^{\text{(Last)}}\}$), where the Pooling function is performed on the set of all the node representations, and it can be Global_mean or Global_max or other more complex pooling functions (Ying et al., 2018; Knyazev et al., 2019).
3 Spectral Analysis of Size-induced Distribution Shifts
In this section, we first show that the independence of the eigenvalue distribution of the propagation matrix ${\mathbf{T}}$ from the graph size is the key to achieving size generalizability of GNNs (§ 3.1). Next, focusing on biologically data, we empirically verify that the eigenvalue distribution of the propagation matrix depends on the graph size (§ 3.2). Finally, we explore the subgraph patterns responsible for the spectral disparities between small and large graphs, unveiling two key findings in § 3.3:
  • Breaking cycles in graphs amplifies the spectrum difference between smaller and larger graphs.
  • Extending cycle lengths in smaller graphs to match larger ones reduces the spectrum difference.
3.1 Graph Spectrum and Size Generalizability of GNNs
To understand how GNNs generalize over graphs with different sizes, we examine the formulation of graph representations. In the context of spectral GNNs, graph representations rely on the eigenvalues of the propagation matrix. Consequently, the connection between graph representations and graph size reduces to the connection between the graph’s spectrum and its size. More formally, we theoretically show the following proposition in Appendix A.
Proposition 3.1.
When graphs of various sizes exhibit distinct eigenvalue distributions for the propagation matrix, the representations learned by spectral GNNs correlate with the graph size.
The proposition suggests that for GNNs to achieve effective generalization to larger graphs, the disparity in the spectrum between small and large graphs should be small.
3.2 Size-related Spectrum Differences in Real Datasets
We now investigate how the eigenvalue distribution of the normalized adjacency matrix varies with graph size in real-world data. As indicated in 3.1, the spectrum discrepancy between small and large graphs affects the size generalizability of GNNs.
Datasets. We explore five pre-processed biological datasets (BBBP, BACE, NCI1, NCI109, and PROTEINS) from the Open Graph Benchmark (Hu et al., 2020b) and TuDataset (Morris et al., 2020). More details about the datasets are provided in Appendix B.
Figure 1: Pairwise graph distance of eigenvalue distributions: Graphs are sorted from small to large, and the (i,j)-th pixel in the plot represents the Wasserstein distance between the i-th and j-th graphs. Dark blue represents a small distance (high similarity), while light red represents a large distance (low similarity). We find that eigenvalue distributions show a strong correlation with the graph size.
Table 1: Average Wasserstein distance between graphs of ‘similar sizes’ and graphs of ‘different sizes’ based on eigenvalue distributions, respectively. The relative difference is computed by the difference of the Wasserstein distance normalized by the Wasserstein distance of similar graphs.
BBBP BACE PROTEINS NCI109 NCI1
Different Size 0.00566 0.00411 0.00765 0.00563 0.00566
Similar Size 0.00184 0.00149 0.00261 0.00215 0.00215
Relative Difference 208% 177% 193% 162% 164%
Figure 2: Pairwise graph distance measured by the Wasserstein distance of eigenvalue distributions after breaking cycles and aligning cycle lengths on the NCI109 dataset. Breaking cycles amplifies the correlation between eigenvalue distribution and graph size, while aligning cycle lengths reduces the correlation.
Table 2: Average Wasserstein distance of eigenvalue distributions between graphs of similar size and graphs of different sizes after breaking cycles and aligning cycle lengths. Relative difference is computed as in Table 1. We use $\uparrow$ ($\downarrow$) to denote the increase (decrease) in the relative difference compared to not taking the corresponding action. Breaking cycles results in a larger relative difference, while aligning cycle lengths reduces the relative difference.
Breaking cycles Aligning cycle lengths
BBBP 0.00545 0.00152 $\uparrow$ 50% 0.00565 0.00211 $\downarrow$ 41%
BACE 0.00420 0.00148 $\uparrow$   6% 0.00417 0.00176 $\downarrow$ 41%
NCI1 0.00547 0.00173 $\uparrow$ 53% 0.00566 0.00242 $\downarrow$ 31%
NCI109 0.00548 0.00174 $\uparrow$ 52% 0.00568 0.00245 $\downarrow$ 31%
PROTEINS 0.00670 0.00212 $\uparrow$ 31% 0.00763 0.00302 $\downarrow$ 41%
Setup. Figure 1 illustrates the pairwise distances of the graphs arranged in ascending order of size, where the distances are calculated using the Wasserstein distance (Villani et al., 2009). We represent the graphs by their empirical distributions of the eigenvalues that are obtained from the normalized adjacency matrix as suggested in (Kipf & Welling, 2017): ${\mathbf{T}}=({\mathbf{D}}+{\mathbf{I}})^{-1/2}({\mathbf{A}}+{\mathbf{I}})({\mathbf{D}}+{\mathbf{I}})^{-1/2}$. Using the normalized Laplacian matrix leads to similar observations. We note that the eigenvalues do not scale with the graph size, and they are bounded between [-1,1]. Dark blue represents a small distance (high similarity) while light red represents a large distance (low similarity).
Results. As can be seen in the three subplots in Figure 1, there is a wide blue band along the diagonal, which indicates that graphs of similar size have more similar eigenvalue distributions than graphs of different sizes. This suggests a strong correlation between the eigenvalue distributions and the graph size. To verify the observation quantitatively, we compute the distance of graphs with ‘similar size’ and graphs of ‘different sizes’ in Table 1. For each graph, we consider the 20 most ‘similar graphs’ in terms of size, and treat the remaining graphs as graphs of ‘different sizes’. The table shows that the Wasserstein distances of eigenvalue distributions between the graphs of different sizes are significantly larger than the distances between graphs of similar size. Based on the empirical results and 3.1, the correlation between the eigenvalue distributions and the graph size results in the correlation of the final graph representation and the graph size, which prevents GNNs from generalizing over larger size.
3.3 Key Findings: Size-related Differences in Subgraph Patterns
In this subsection, we aim to identify the subgraph patterns that explain the spectrum differences between small and large graphs. Our empirical analysis pinpointed several peaks in the graph spectrum that match the spectrum of cycles. This motivated us to examine how cycle properties differ in small and large graphs and how these differences are revealed in the spectrum. Specifically, we aim to answer two questions: (Q1) How does the existence of cycles in the graphs influence the spectrum differences? (Q2) How do the variations in cycle lengths contribute to differences in the spectrum? In our analysis, we investigate the properties of the cycles in the cycle basis of each graph.
3.3.1 Existence of Cycles & Spectrum: The Impact of Breaking Cycles
To understand how the existence of cycles in the graphs influences the spectrum differences, we break all basis cycles with minimal edge removal while maintaining the same number of disconnected components, according to the details and algorithm given in Section E.1. We analyze the impact of breaking cycles by assessing the corresponding changes in the spectrum. By following the convention of Section 3.2, in the center of Figure 2, we plot the pairwise graph distance based on eigenvalue distributions of graphs with different sizes after breaking cycles. The blue band along the diagonal of the plot becomes darker and narrower, suggesting a larger spectrum difference between small and large graphs and a stronger correlation between the spectrum and graph size. To evaluate the effects quantitatively, we further compute the changes in the relative spectrum difference and present the results in Table 2. These results indicate that failing to consider cycle information can lead to more significant differences in the spectrum between graphs of varying sizes, potentially causing GNNs to struggle with generalizing effectively to larger graphs.
3.3.2 Cycle Length & Spectrum: Aligning Cycle Lengths
In Section 3.3.1, we showed that cycle information is crucial for GNNs to achieve size generalizability. We now further explore what cycle information helps reduce the spectrum difference between small and large graphs. To facilitate our exploration, we divide each real-world dataset into two subsets: one subset contains small graphs, and the other subset containts graphs of significantly larger size. Further details regarding this dataset split can be found in Appendix B. Using this dataset split, we observe a significant difference in the cycle lengths for small and large graphs (Section E.1). As described in Section E.1, to reduce that difference, we align the average cycle lengths between small and large graphs by randomly inserting redundant nodes to increase the cycle lengths in small graphs. The rightmost heatmap in Figure 2 shows how the correlation of eigenvalue distributions and graph size changes after aligning cycle lengths. We observe a lighter blue band along the diagonal, which suggests a weaker correlation between the spectrum and graph size. Furthermore, Table 2 quantitatively presents the changes in the relative spectrum difference between small and large graphs. We observe that aligning cycle lengths results in reduced disparities in the spectrum between graphs of different sizes. This indicates that GNNs capable of generalizing across varying cycle lengths may exhibit better size generalizability. In Section E.2, we show that our approach of aligning the cycle lengths is more effective at reducing the spectrum disparities than randomly adding the same number of nodes and edges. Importantly, our observation on cycle length variation differs from the assumptions in (Bevilacqua et al., 2021). While they assume the density invariance of the induced k-sized subgraphs, we observe the absence of large cycles in small graphs, indicating a lack of density invariance for cycles.
4 Methodology: Size-insensitive Attention
Our findings in Section 3 suggest that GNNs with better ability to identify cycles and generalize over cycle lengths may have better size generalizability on biological graphs. However, recent work (Chen et al., 2020) has found that most GNNs are incapable of learning cycle information. Inspired by these, we propose to incorporate cycle information into GNNs by encoding it in the features and leveraging them within the attention mechanism to guide the learning process. Specifically, for each graph $\mathcal{G}$, we obtain its cycle basis $\mathcal{C}$. Then, for each node $v_{i}\in\mathcal{G}$, we calculate the average length of the cycle basis to which it belongs:
$\ell_{i}=\begin{cases}\frac{\sum_{j=1}^{|\mathcal{C}|}|\mathcal{C}_{j}|\cdot\mathbbm{1}_{\{v_{i}\in\mathcal{C}_{j}\}}}{\sum_{j=1}^{|\mathcal{C}|}\mathbbm{1}_{\{v_{i}\in\mathcal{C}_{j}\}}}&\text{ if }v_{i}\text{ belongs to some cycles}\\ 0&\text{ otherwise,}\end{cases}$
$\mathbf{c}_{i}=[\mathbbm{1}_{\{v_{i}\in\text{cycle}\}},\ell_{i}].$
${\mathbf{k}}=\texttt{Softmax}(\mathbf{C}{\mathbf{w}}^{\top}_{A})\cdot N,\;\;\;\mathbf{x}^{\mathcal{G}}=\texttt{Global\_max}(\texttt{Diag}({\mathbf{k}})\cdot\mathbf{X}^{(\text{Last})})$
where ${\mathbf{w}}^{\top}_{A}$ is a learnable vector, and $\texttt{Diag}(\cdot)$ creates a diagonal matrix using the vector as its elements. We note that when we train on small graphs and test on large graphs, some structural features may not be seen in the training, such as certain cycle lengths in the large graphs. We rely on the attention mechanism to generalize to those cases. Though simple as it seems, SIA is found more effective in size generalization than expressive GNNs which are good at identifying cycles Table 3.
5 Experiments
In this section, we conduct extensive experiments to evaluate our proposed strategy SIA. We aim to answer the following research questions: (RQ1) Is cycle information helpful for size generalization?: Do strategies incorporating cycle information effectively enhance the size generalizability of GNNs? (RQ2) Comparison with other baselines: How does SIA compare to other baseline strategies?
5.1 Setup
Dataset. We use the same biological datasets as in Section 3.2. More details can be found in Appendix B.
Data Preprocessing and Important Training Details. In order to analyze size generalizability, we have four splits for each dataset: train, validation, small_test, and large_test, where large_test contains graphs with significantly larger sizes. We generate the splits as follows. First, we sort the samples in the dataset by their size. Next, We take the train, validation, and small_test split from the 50% smallest graphs in the dataset. An intuitive way of getting the large_split is to take the top k largest graphs. However, doing so would result in severe label skewness (class imbalances) between the small_test and the large_test as demonstrated by Table 5 in Appendix C. To avoid such a severe label shift, we select the same number of graphs per class as in the small_test subset, starting from the largest graph within each class. This way guarantees that the label distribution between small_test and large_test is the same, while ensuring that the graph size in the latter is 2-10 times larger. Nevertheless, the smallest 50% samples still have significant class imbalance. To address this issue, we use upsampling during training throughout the experiments, and we use F1 as the metric to measure the model performance. More details about data preprocessing, hyperparameters, and training can be found in Appendix C and Appendix G.
Baselines. We use six neural network models as our GNN backbones. Each model consists of three layers, with a global max pooling layer employed in the final layer. The baseline models are: Multilayer Perceptron (MLP), GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2018), GIN (Xu et al., 2018), FAGCN (Bo et al., 2021), and GNNML3 (Balcilar et al., 2021b). We integrate seven model-agnostic strategies with these GNN backbones. We use (1) +SIA to denote our proposed size-insensitive attention method. To further validate that cycle information is helpful for size generalization, we propose two additional baselines: (2) +SSL introduces a self-supervised loss in addition to the standard classification loss, where the auxiliary task is to predict whether a node is part of a cycle; and (3) +AugCyc reduces the discrepancy between small and large graphs by augmenting the training set through the extension of cycle lengths. Further details about +SSL and +AugCyc can be found in Appendix D. We also compare with other model-agnostic strategies: (4) SizeShiftReg (Buffelli et al., 2022) (+SSR), a regularization based on the idea of simulating a shift in the size of the training graphs using coarsening techniques; (5) RPGNN (Murphy et al., 2019) (+RPGNN), an expressive model for arbitrary-sized graphs; two versions of CIGA ((6)+CIGAv1 & (7)+CIGAv2(Chen et al., 2022), a causal model good at handling out-of-distribution problems on graphs. Besides these model-agnostic strategies, our baselines also include an expressive model (8) SMP (Vignac et al., 2020) (SMP), which excels at the cycle detection task.
Table 3: Size generalizability evaluated by the graph classification performance on small and large test graphs. The performance is reported by the average F1 scores and its standard deviation. The rightmost column denotes the average improvements compared with the original performance using the same backbone model across five different datasets. The largest average improvement within the same model and small/large category is highlighted in orange. All strategies enhance GNNs’ size generalizability, with +SIA surprisingly emerging as the most effective method.

### v5

# Tackling Size Generalization of Graph Neural Networks on Biological Data from a Spectral Perspective

@@ 4.1. Strategy S1: Size-insensitive Attention (SIA)
4.2. Strategy S2: Self-supervised Auxiliary Task
Our second proposed strategy utilizes a self-supervised (SSL) auxiliary task to enhance the node representations with cycle-related information. The auxiliary task is to predict whether a node belongs to a cycle. We do not utilize cycle lengths as labels because large test graphs may have cycle lengths not present in the training data. Formally, let $X^{(\text{Last})}$ denote the node representations obtained after the last graph convolution. The conventional way of learning is to directly apply a pooling operator and then minimize the loss function for label supervision as below:
(5) $\mathcal{L}_{\text{label}}=\texttt{CrossEntropy}(\texttt{Linear}(\texttt{Pooling}({\mathbf{X}}^{(\text{Last})}),y^{{\mathcal{G}}})),$
(6) $\mathcal{L}_{\text{cycle}}=\texttt{CrossEntropy}(\texttt{MLP}({\mathbf{X}}^{(\text{Last})}),{\mathbf{y}}_{\text{cycle}}),$
(7) $\mathcal{L}=\mathcal{L}_{\text{label}}+\lambda\mathcal{L}_{\text{cycle}},$
where $\lambda$ is a hyperparameter tuned via cross-validation.
4.3. Strategy S3: Augmentation
Our third strategy, +AugCyc, aims to reduce the discrepancy between small and large graphs through direct augmentation for the small training graphs. We augment the training graphs by extending the cycle lengths such that the average cycle length and standard deviation align with those in large graphs. To achieve this augmentation, we use the algorithm detailed in Section G.1. In short, our algorithm selectively increases the cycle length of a subset of small graphs. The newly added nodes share the same features from the nodes with the lowest degrees in the same cycle. Last, we feed the augmented training graphs to GNNs for graph classification.
5. Experiments
In this section, we conduct extensive experiments to evaluate our proposed strategies. Starting with real datasets, we aim to answer the following research questions:
$\bullet$ (RQ1) Is cycle information helpful for size generalization? Do our strategies incorporating cycle information effectively enhance the size generalizability of GNNs?
$\bullet$ (RQ2) How do our strategies compare to other baseline strategies?
$\bullet$ (RQ3) How effective is SIA on large datasets?
$\bullet$ (RQ4) How effective is it on synthetic datasets that simulate real-world molecular graphs with varying cycle distributions. [^1] The code is available at https://github.com/GaotangLi/SizeGenBio.
5.1. Experimental Setup
Real Datasets. We use the biological datasets in Section 3.1.
Synthetic Datasets. As discussed in Section 3.2.3, cyclic structures are fundamental in chemical molecules. We consider a binary classification problem: the first category consists of cycles and connections without shared bonds, while the second includes fused cycles (two or more cycles sharing bonds) with additional connections. Each graph is connected, and all features are assigned as i.i.d. nine-dimensional Gaussian vectors, aligning with the feature dimensionality of BBBP. Our dataset reflects real-world molecular structures, where cycle fusion significantly alters molecular properties. Fused rings affect physical characteristics, chemical reactivity, and biological activity. For example, benzene ($C_{6}H_{6}$) is a highly stable six-membered aromatic ring, whereas naphthalene ($C_{10}H_{8}$), composed of two fused benzene rings, exhibits increased reactivity in electrophilic substitution and higher melting and boiling points. Similarly, while cyclohexane ($C_{6}H_{12}$) is chemically inert, decalin ($C_{10}H_{18}$), formed by fusing two cyclohexane rings, becomes significantly more rigid.
Data Preprocessing and Important Training Details. In order to analyze size generalizability, we generate four splits for each dataset: train, validation, small_test, and large_test, where large_test contains graphs with significantly larger sizes. We generate the splits as follows. First, we sort the samples in the dataset by their size. Next, we take the train, validation, and small_test split from the 50% smallest graphs in the dataset. An intuitive way of getting the large_split is to take the top-k largest graphs. However, doing so would result in severe label skewness (class imbalances) between the small_test and the large_test. To avoid such a severe label shift, we select the same number of graphs per class as in the small_test subset, starting from the largest graph within each class. This way guarantees that the label distribution between small_test and large_test is the same, while ensuring that the graph size in the latter is 2-10 times larger. Nevertheless, the smallest 50% samples still have significant class imbalance. To address this issue, we use upsampling during training throughout the experiments, and we use F1 as the metric to measure the model performance. More details about data preprocessing, hyperparameters, and training can be found in Appendix C and Appendix D.
Baselines. We consider a total of 43 baselines: a strong expressive model, SMP (Vignac et al., 2020), which excels at the cycle detection task, and the combination of 6 backbone models with 7 model-agnostic strategies. Specifically, the baseline models are: Multilayer Perceptron (MLP), GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GIN (Xu et al., 2018), FAGCN (Bo et al., 2021), and GNNML3 (Balcilar et al., 2021b). We integrate seven model-agnostic strategies with these GNN backbones: our three proposed strategies, (1) +SIA, size-insensitive attention method that leverages the attention mechanism to incorporate cycle information (S1); (2) +SSL, our self-supervised based strategy that encourages the node representation to contain cycle information (S2); (3) +AugCyc, our augmentation-based strategy that aligns cycle lengths (S3); as well as other existing model-agnostic strategies: (4) (4) SizeShiftReg (Buffelli et al., 2022) (+SSR), a regularization based on the idea of simulating a shift in the size of the training graphs using coarsening techniques; (5) RPGNN (Murphy et al., 2019) (+RPGNN), an expressive model for arbitrary-sized graphs; and (6)+CIGAv1 & (7)+CIGAv2 (Chen et al., 2022), two versions of a causal model that effectively handles out-of-distribution problems on graphs.
Table 3. Size generalizability evaluated by the graph classification performance on small and large test graphs. The performance is reported by the average F1 scores and its standard deviation. The rightmost column denotes the average improvements compared with the original performance using the same backbone model across five different datasets. The largest average improvement within the same model and small/large category is highlighted in orange. All strategies enhance GNNs’ size generalizability, with +SIA drop as the most effective method.